OpenStudy (anonymous):

Jason wants to choose 9 players for his track team. There are 12 players to choose from. How many different teams can Jason make

OpenStudy (anonymous):

nCr=n!/(n-r)!r! 12C9=12!/\[\frac{ 12! }{ ? }\](12-9)!9! =12!/3!9! =12x11x10x9x8x7x6x5x4x3x2/3x2x9x8x7x6x5x4x3x2 cancel and solve

OpenStudy (anonymous):

@CiroR whats the answer then?